Find a two-decimal-place number between 0 and one such that the sum of its digits is 9 and such that if the digits are reversed the number is increased by 0.27
Answer by robertb(5830)

The number is 0.36.
Let x be the tens place, y the ones place.
Then x + y = 9, and 10x + y + 27 = 10y + x
==> 27 = 9y - 9x <==> y - x = 3
y - x = 3
y + x = 9
==> 2y = 12
y = 6 ==> x = 3.
The decimal number is 0.36.